# Yang-Baxter Hochschild Cohomology

### v3

@@ Abstract.
Braided algebras are associative algebras endowed with a Yang-Baxter operator that satisfies certain compatibility conditions involving the multiplication. Along with Hochschild cohomology of algebras, there is also a notion of Yang-Baxter cohomology, which is associated to any Yang-Baxter operator. In this article, we introduce and study a cohomology theory for braided algebras in dimensions 2 and 3, that unifies Hochschild and Yang-Baxter cohomology theories. We show that its second cohomology group classifies infinitesimal deformations of braided algebras. We provide infinite families of examples of braided algebras, including Hopf algebras, tensorized multiple conjugation quandles, and braided Frobenius algebras. Moreover, we derive the obstructions to quadratic deformations, and show that these obstructions lie in the third cohomology group. Relations to Hopf algebra cohomology are also discussed.
@@ 2.1. Braided algebras from tensorized multiple conjugation quandles
First, recall a quandle [Joyce82, Matveev82], is a non-empty set $X$ with a binary operation *$*:X\times X\to X$ satisfying the following axioms.
@@
A multiple conjugation quandle (MCQ) $X$ is the disjoint union of groups $G_{\lambda}$, where $\lambda$ is an element of an index set $\Lambda$, with a binary operation *$*:X\times X\to X$ satisfying the following axioms.
@@
Let $X=\sqcup_{\lambda\in\Lambda}G_{\lambda}$ be an MCQ with a quandle operation ***. Let ${\mathbb{k}}[X]$ be the free ${\mathbb{k}}$-module generated by $X$. Define a multiplication $\mu:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]$ on generators by $\mu(x\otimes y):=xy$ if $x,y\in G_{\lambda}$ for some $\lambda\in\Lambda$, and $\mu(x\otimes y):=0$ otherwise, and extended linearly. Define $R:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]$ on generators $x,y\in X$ by $R(x\otimes y):=y\otimes(x*y)$ and extended linearly. Then ${\mathbb{k}}[X]$ is a braided algebra, as we now proceed to show. If two of $x,y,z\in X$ belong to distinct $G_{\lambda}$’s, then the triple product is 00, and if all belong to the same $G_{\lambda}$ then it is associative, hence $\mu$ is associative. Equations (1) and (2), respectively, follow from the conditions $(ab)*x=(a*x)(b*x)$ and $x*(ab)=(x*a)*b$ for any $x\in X$, $a,b\in G_{\lambda}$ on generators, and if $a,b$ belong to distinct $G_{\lambda}$ then the values are 00 for both sides of the equations, hence ${\mathbb{k}}[X]$ is a YI and IY algebra.
@@ 3.2. Yang-Baxter cohomology up to dimension 3
Deformation cohomology theory was developed in [Eisermann] for all dimensions. The cochains in [Eisermann] differ from ours by compositions of YB operators. For example, the 2-cochain $c^{\prime}$ in [Eisermann] is written as $Rc$ for our 2-cochain $c$. We adopt our definition and diagrammatics along the line of [CCES-coalgebra, CCES-adjoint] for the purpose of unifying Yang-Baxter cohomology with algebra cohomology.
@@ 3.4. Infinitesimal deformations of braided algebras
Recall that in order for $\tilde{\mu}$ to satisfy the associativity condition, $\psi$ needs to be a Hochschild $2$-cocycle, while $\tilde{R}$ is a YB operator if and only if $\phi$ is a YB $2$-cocycle. Since the latter fact is not found as commonly in the literature as the former, we give a brief proof of it. See also [Eisermann, Eisermann1]. We set $\phi_{0}:=R$ and $\phi_{1}:=\phi$ for notational convenience. The YBE for $\tilde{R}$ then gives us for the left hand side (modulo terms of quadratic order or higher in $\hbar$)
@@
To prove the theorem, we need to show two facts. First, that Yang-Baxter Hochschild $2$-cocycles define infinitesimal deformations, and that cobounded $2$-cocycles produce equivalent deformations. Second, that each infinitesimal deformation arises from a $2$-cocycle, and that if two deformations are isomorphic, then the corresponding $2$-cocycles are cobounded in Yang-Baxter Hochschild cohomology. The first part, that a Yang-Baxter Hochschild $2$-cocycle gives a Yang-Baxter Hochschild infinitesimal deformation, follows from Lemma 3.6. Suppose that $(\phi,\psi)$ is cobounded in the Yang-Baxter Hochschild cohomology, where $\phi:V\otimes V\rightarrow V\otimes V$ and $\psi:V\otimes V\rightarrow V$. We show that the induced infinitesimal deformation is equivalent to the original. By definition of first differential, the assumption means that $\phi=\delta_{\rm YB}f$ and $\psi=\delta_{\rm H}f$ for some $f:V\rightarrow V$. Then, $\tilde{R}$ is the trivial YB deformation and $\tilde{\mu}$ is the trivial algebra deformation. Then, we can construct the map $\tilde{f}=\mathbb{1}+\hbar f$. This is invertible in $V\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$ with inverse $\tilde{f}^{-1}=\mathbb{1}-\hbar f$. Also, we now show that both $\tilde{f}$ and $\tilde{f}^{-1}$ are braided algebra homomorphisms between the undeformed braided algebra $(V,\mu,R)$ and the deformed braided algebra $(\tilde{V},\tilde{\mu},\tilde{R})$. We show that $(\tilde{f}\otimes\tilde{f})R=\tilde{R}(\tilde{f}\otimes\tilde{f})$. For the left hand side of the previous equation we have, up to terms of higher order in $\hbar$,
@@
The two terms are seen to be equal since $\delta^{1}_{\rm YB}f=\phi$. A similar inspection of the homomorphism condition for associative algebras shows also that $\tilde{f}\tilde{\mu}=\mu(\tilde{f}\otimes\tilde{f})$. A similar procedure shows that the inverse of $\tilde{f}$ is a braided algebra homomorphism, completing the first part of the proof. The statement that if $\tilde{R}$ and $\tilde{\mu}$ produce an infinitesimal deformation of braided algebras, then $(\phi,\psi)$ is a Yang-Baxter Hochschild $2$-cocycle, was shown in Lemma 3.6. To complete the proof, we only need to show that if $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$ give rise to equivalent deformations, then they are cohomologous. Let $\tilde{f}$ be a map that gives the isomorphism between the braided algebra structures associated to $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$. Then, since $\tilde{f}$ is an isomorphism that fixes the degree zero braided algebra structure, it follows that $\tilde{f}=\mathbb{1}+\hbar f$ for some $f$. Now, up to degrees higher than $1$ in $\hbar$ we have that $\tilde{f}^{-1}={\mathbb{1}}-\hbar f$. Since $\tilde{f}$ is a homomorphism of braided algebra structures it follows that
@@ 5. Hopf algebra cohomology and braided algebra cohomology
Therefore, we can rewrite $S^{\prime}$ using these symmetries, and proceed in a way analogous to the previous result to obtain that $\xi(S\otimes\mathbb{1})\Delta+\mu(S\otimes\mathbb{1})\zeta+\mu(S^{\prime}\otimes\mathbb{1})\Delta=0$ as well, which completely shows that $S^{\prime}$ satisfies Equation (5). The fact that $S^{\prime}$ is unique, follows from uniqueness of antipodes in Hopf algebras. We will denote the normalized second cohomology group of $H$ with coefficients in $H$ with the symbol $\hat{H}^{2}_{\rm Hf}(H,H):=\hat{Z}_{\rm Hf}^{2}(H,H)/(B^{1}_{\rm Hf}(H,H)\cap\hat{Z}_{\rm Hf}^{2}(H,H))$.
@@ 6.1. The third differentials
$\delta^{3,2}_{\rm YI}(\alpha\oplus\alpha^{\prime})$ $=$ $(\mu\otimes{\mathbb{1}}^{\otimes 2})\alpha^{\prime}+(\alpha\otimes{\mathbb{1}})({\mathbb{1}}\otimes R)((\mu\otimes{\mathbb{1}})+({\mathbb{1}}\otimes R\otimes{\mathbb{1}})(R\otimes{\mathbb{1}}^{\otimes 2}))({\mathbb{1}}\otimes\alpha)(\mu\otimes{\mathbb{1}})$
$-({\mathbb{1}}^{\otimes 2}\otimes\mu)\alpha-({\mathbb{1}}\otimes\alpha^{\prime})(R\otimes{\mathbb{1}})({\mathbb{1}}\otimes\mu)-({\mathbb{1}}\otimes R\otimes{\mathbb{1}})({\mathbb{1}}^{\otimes 2}\otimes R)(\alpha^{\prime}\otimes{\mathbb{1}})({\mathbb{1}}\otimes\mu).$
@@
$\delta^{3,3}_{\rm YI}(\alpha\oplus\gamma)$ $=$ $(\gamma\otimes{\mathbb{1}})R+({\mathbb{1}}\otimes\mu)\alpha+({\mathbb{1}}\otimes\alpha)(R\otimes{\mathbb{1}})\mu$
$-(\mu\otimes{\mathbb{1}})\alpha-({\mathbb{1}}^{\otimes 2}\otimes R)(\alpha\otimes{\mathbb{1}}^{\otimes 2})\mu-({\mathbb{1}}^{\otimes 2}\otimes R)({\mathbb{1}}\otimes R\otimes{\mathbb{1}})(R\otimes{\mathbb{1}}^{\otimes 2})\gamma.$
@@ 6.2. Quadratic deformation for Hochschild cohomology
where $\Theta_{k+1}$ contains terms not involving $\mu_{k+1}$. Similarly as above, it holds that $\delta^{3}_{H}(\Theta_{k+1})=0$, and the obstruction to extend a degree $k$ deformation to a degree $k+1$ deformation lies in the third cohomology group [Gerst]. In this case we do not provide a diagrammatic proof of these classic results, as the analogue approach for braided algebras becomes increasingly complicated to handle. However, we expect that the results below that show that the obstruction for degree $2$ deformations lie in the third cohomology group of braided algebras, would extend to the case of higher degree deformations as in the case of associative algebras.
@@ 6.3. Higher order deformations for Yang-Baxter Hochschild homology
We investigate the obstructions to higher deformations for braided algebras, and obtain integrability criteria at least in some special cases of interest. This requires additional obstructions for the compatibility among higher degree deformation terms lying in the third YBH cohomology groups. First, we pose the following inductive definition, where Defintion 3.4 serves as the base case $n=1$.

### v4

@@ Abstract.
Braided algebras are associative algebras endowed with a Yang-Baxter operator that satisfies certain compatibility conditions involving the multiplication. Along with Hochschild cohomology of algebras, there is also a notion of Yang-Baxter cohomology, which is associated to any Yang-Baxter operator. In this article, we introduce and study a cohomology theory for braided algebras in dimensions 2 and 3, that unifies Hochschild and Yang-Baxter cohomology theories, and generalizes to all dimensions in characteristic $2$. We show that its second cohomology group classifies infinitesimal deformations of braided algebras. We provide infinite families of examples of braided algebras, including Hopf algebras, tensorized multiple conjugation quandles, and braided Frobenius algebras. Moreover, we derive the obstructions to higher deformations, which lie in the third cohomology group. Relations to Hopf algebra cohomology are also discussed.
@@ 2.1. Braided algebras from tensorized multiple conjugation quandles
First, recall a quandle [Joyce82, Matveev82], is a non-empty set $X$ with a binary operation $*:X\times X\to X$ satisfying the following axioms.
@@
A multiple conjugation quandle (MCQ) $X$ is the disjoint union of groups $G_{\lambda}$, where $\lambda$ is an element of an index set $\Lambda$, with a binary operation $*:X\times X\to X$ satisfying the following axioms.
@@
Let $X=\sqcup_{\lambda\in\Lambda}G_{\lambda}$ be an MCQ with a quandle operation *. Let ${\mathbb{k}}[X]$ be the free ${\mathbb{k}}$-module generated by $X$. Define a multiplication $\mu:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]$ on generators by $\mu(x\otimes y):=xy$ if $x,y\in G_{\lambda}$ for some $\lambda\in\Lambda$, and $\mu(x\otimes y):=0$ otherwise, and extended linearly. Define $R:{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]\rightarrow{\mathbb{k}}[X]\otimes{\mathbb{k}}[X]$ on generators $x,y\in X$ by $R(x\otimes y):=y\otimes(x*y)$ and extended linearly. Then ${\mathbb{k}}[X]$ is a braided algebra, as we now proceed to show. If two of $x,y,z\in X$ belong to distinct $G_{\lambda}$’s, then the triple product is 00, and if all belong to the same $G_{\lambda}$ then it is associative, hence $\mu$ is associative. Equations (1) and (2), respectively, follow from the conditions $(ab)*x=(a*x)(b*x)$ and $x*(ab)=(x*a)*b$ for any $x\in X$, $a,b\in G_{\lambda}$ on generators, and if $a,b$ belong to distinct $G_{\lambda}$ then the values are 00 for both sides of the equations, hence ${\mathbb{k}}[X]$ is a YI and IY algebra.
@@ 3.2. Yang-Baxter cohomology up to dimension 3
A deformation cohomology theory for Yang-Baxter operators was developed further to all higher dimensions in [SZ-BC] extending those in this paper, using the diagrammatic method developed above.
This YB cohomology theory is different from those in [Eisermann]; the differentials in [Eisermann] differ from ours by compositions of YB operators. For example, the 2-cochain $c^{\prime}$ in [Eisermann] is written as $Rc$ for our 2-cochain $c$. We adopt our definition and diagrammatics along the line of [CCES-coalgebra, CCES-adjoint] for the purpose of unifying Yang-Baxter cohomology with algebra cohomology.
@@ 3.3. Yang-Baxter Hochschild cohomology up to dimension $3$
Remark 3.4.
The YBH cohomology is further generalized in [SZ-BC] to all higher dimensions in a multicomplex form, that extends the definition for low dimensions presented in this paper.
@@ 3.4. Infinitesimal deformations of braided algebras
Recall that in order for $\tilde{\mu}$ to satisfy the associativity condition, $\psi$ needs to be a Hochschild $2$-cocycle, while $\tilde{R}$ is a YB operator if and only if $\phi$ is a YB $2$-cocycle.
Since the latter fact is not found as commonly in the literature as the former, we give a brief proof of it. See also [Eisermann, Eisermann1]. We set $\phi_{0}:=R$ and $\phi_{1}:=\phi$ for notational convenience. The YBE for $\tilde{R}$ then gives us for the left hand side (modulo terms of quadratic order or higher in $\hbar$)
@@
To prove the theorem, we need to show two facts. First, that Yang-Baxter Hochschild $2$-cocycles define infinitesimal deformations, and that cobounded $2$-cocycles produce equivalent deformations. Second, that each infinitesimal deformation arises from a $2$-cocycle, and that if two deformations are isomorphic, then the corresponding $2$-cocycles are cobounded in Yang-Baxter Hochschild cohomology. The first part, that a Yang-Baxter Hochschild $2$-cocycle gives a Yang-Baxter Hochschild infinitesimal deformation, follows from Lemma 3.7.
Suppose that $(\phi,\psi)$ is cobounded in the Yang-Baxter Hochschild cohomology, where $\phi:V\otimes V\rightarrow V\otimes V$ and $\psi:V\otimes V\rightarrow V$. We show that the induced infinitesimal deformation is equivalent to the original. By definition of first differential, the assumption means that $\phi=\delta_{\rm YB}f$ and $\psi=\delta_{\rm H}f$ for some $f:V\rightarrow V$. Then, $\tilde{R}$ is the trivial YB deformation and $\tilde{\mu}$ is the trivial algebra deformation. Then, we can construct the map $\tilde{f}=\mathbb{1}+\hbar f$. This is invertible in $V\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$ with inverse $\tilde{f}^{-1}=\mathbb{1}-\hbar f$. Also, we now show that both $\tilde{f}$ and $\tilde{f}^{-1}$ are braided algebra homomorphisms between the undeformed braided algebra $(V,\mu,R)$ and the deformed braided algebra $(\tilde{V},\tilde{\mu},\tilde{R})$. We show that $(\tilde{f}\otimes\tilde{f})R=\tilde{R}(\tilde{f}\otimes\tilde{f})$. For the left hand side of the previous equation we have, up to terms of higher order in $\hbar$,
@@
The two terms are seen to be equal since $\delta^{1}_{\rm YB}f=\phi$. A similar inspection of the homomorphism condition for associative algebras shows also that $\tilde{f}\tilde{\mu}=\mu(\tilde{f}\otimes\tilde{f})$. A similar procedure shows that the inverse of $\tilde{f}$ is a braided algebra homomorphism, completing the first part of the proof.
The statement that if $\tilde{R}$ and $\tilde{\mu}$ produce an infinitesimal deformation of braided algebras, then $(\phi,\psi)$ is a Yang-Baxter Hochschild $2$-cocycle, was shown in Lemma 3.7.
To complete the proof, we only need to show that if $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$ give rise to equivalent deformations, then they are cohomologous. Let $\tilde{f}$ be a map that gives the isomorphism between the braided algebra structures associated to $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$. Then, since $\tilde{f}$ is an isomorphism that fixes the degree zero braided algebra structure, it follows that $\tilde{f}=\mathbb{1}+\hbar f$ for some $f$. Now, up to degrees higher than $1$ in $\hbar$ we have that $\tilde{f}^{-1}={\mathbb{1}}-\hbar f$. Since $\tilde{f}$ is a homomorphism of braided algebra structures it follows that
@@ 5. Hopf algebra cohomology and braided algebra cohomology
Therefore, we can rewrite $S^{\prime}$ using these symmetries, and proceed in a way analogous to the previous result to obtain that $\xi(S\otimes\mathbb{1})\Delta+\mu(S\otimes\mathbb{1})\zeta+\mu(S^{\prime}\otimes\mathbb{1})\Delta=0$ as well, which completely shows that $S^{\prime}$ satisfies Equation (5). The fact that $S^{\prime}$ is unique, follows from uniqueness of antipodes in Hopf algebras.
We will denote the normalized second cohomology group of $H$ with coefficients in $H$ with the symbol $\hat{H}^{2}_{\rm Hf}(H,H):=\hat{Z}_{\rm Hf}^{2}(H,H)/(B^{1}_{\rm Hf}(H,H)\cap\hat{Z}_{\rm Hf}^{2}(H,H))$.
@@ 6.1. The third differentials
$\delta^{3,2}_{\rm YI}(\alpha\oplus\alpha^{\prime})$ $=$ $\alpha^{\prime}(\mu\otimes{\mathbb{1}}^{\otimes 2})+(\mu\otimes{\mathbb{1}})({\mathbb{1}}\otimes R)(\alpha\otimes{\mathbb{1}})+({\mathbb{1}}\otimes\alpha)(\mu\otimes{\mathbb{1}})(R\otimes{\mathbb{1}}^{\otimes 2})({\mathbb{1}}\otimes R\otimes{\mathbb{1}})$
$-\alpha({\mathbb{1}}^{\otimes 2}\otimes\mu)-({\mathbb{1}}\otimes\mu)(R\otimes{\mathbb{1}})({\mathbb{1}}\otimes\alpha^{\prime})-(\alpha^{\prime}\otimes{\mathbb{1}})({\mathbb{1}}\otimes\mu)({\mathbb{1}}^{\otimes 2}\otimes R)({\mathbb{1}}\otimes R\otimes{\mathbb{1}}).$
@@
$\delta^{3,3}_{\rm YI}(\alpha\oplus\gamma)$ $=$ $R(\gamma\otimes{\mathbb{1}})+\alpha({\mathbb{1}}\otimes\mu\otimes{\mathbb{1}})+({\mathbb{1}}\otimes\mu)(R\otimes{\mathbb{1}})({\mathbb{1}}\otimes\alpha)$
$-\alpha(\mu\otimes{\mathbb{1}}^{\otimes 2})-({\mathbb{1}}\otimes\mu)(\alpha\otimes{\mathbb{1}})({\mathbb{1}}^{\otimes 2}\otimes R)-({\mathbb{1}}\otimes\gamma)(R\otimes{\mathbb{1}}^{\otimes 2})({\mathbb{1}}\otimes R\otimes{\mathbb{1}})({\mathbb{1}}^{\otimes 2}\otimes R).$
@@ 6.2. Quadratic deformation for Hochschild cohomology
where $\Theta_{k+1}$ contains terms not involving $\mu_{k+1}$. Similarly as above, it holds that $\delta^{3}_{H}(\Theta_{k+1})=0$, and the obstruction to extend a degree $k$ deformation to a degree $k+1$ deformation lies in the third cohomology group [Gerst].
In this case we do not provide a diagrammatic proof of these classic results, as the analogue approach for braided algebras becomes increasingly complicated to handle. However, we expect that the results below that show that the obstruction for degree $2$ deformations lie in the third cohomology group of braided algebras, would extend to the case of higher degree deformations as in the case of associative algebras.
@@ 6.3. Higher order deformations for Yang-Baxter Hochschild homology
We investigate the obstructions to higher deformations for braided algebras, and obtain integrability criteria at least in some special cases of interest. This requires additional obstructions for the compatibility among higher degree deformation terms lying in the third YBH cohomology groups. First, we pose the following inductive definition, where Defintion 3.5 serves as the base case $n=1$.
@@
Lemma 6.6 is generalized to the extension of any degree $n$ deformation to a degree $n+1$ deformation by considering Markl-Stasheff calculus of deviations.
Theorem 6.8.
Let $\bar{\psi}=\sum_{i=0}^{n}\hbar^{i}\psi_{i}$ and $\bar{\phi}=\sum_{i=0}^{n}\hbar^{i}\phi_{i}$ be a degree $n$ deformation of $\mu=\psi_{0}$ and $R=\phi_{0}$. Let $\psi_{n+1}$, $\phi_{n+1}$, $\Theta_{n+1}$, $\Xi^{\bullet}_{n+1}$, $\Omega^{\bullet}_{n+1}$, $\Lambda^{\bullet}_{n+1}$ be as in Lemma 6.5. Then, we have
$\delta^{3}_{\rm YBH}(\Theta_{n+1}\oplus(\Xi^{\rm IY}_{n+1}-\Omega^{\rm IY}_{n+1})\oplus(\Xi^{\rm YI}_{n+1}-\Omega^{\rm YI}_{n+1})\oplus\Lambda_{n+1})=0.$
In other words, the obstruction to extending a degree $n$ braided algebra deformation to a degree $n+1$ deformation lies in the third YBH cohomology group.
The proof of this result is based on the use of the diagrams in Figure 12 and Figure 14, as done explicitly in [SZ-BC]. In fact, the figures can be used to compute the deviations for higher degrees to hold, from which we can infer that the obstruction to extendind a deformation to a higher degree holds up to the third differentials. Details can be found in [SZ-BC].
Then, Corollary 6.7 is immediately further generalized as follows.
Corollary 6.9.
If $H^{3}_{\rm YBH}(V,V)=0$, then any infinitesimal deformation of $V$ can be extended to a full deformation of arbitrary degree.
We include the proof of Lemma 6.6 since this proof is more diagrammatic and different from methods used in [SZ-BC]. Moreover, the graphs used in the present proof were used in [SZ-BC] in a different manner.